# Variational Source-Channel Coding for Semantic Communication

### v1

@@ Abstract
The rapid development of artificial intelligence (AI) brings new opportunities and challenges to classical communication. AI applications not only impose significant transmission pressure on current communication networks but also affect the way information is stored and utilized due to its potent nonlinear fitting capabilities. Consequently, semantic communication technology emerges as a pivotal bridge connecting AI with classical communication. The current semantic communication systems are generally modeled as an Auto-Encoder (AE). AE lacks a deep integration of AI principles with communication strategies due to its inability to effectively capture channel dynamics. This gap makes it difficult to justify the need for joint source-channel coding (JSCC) and to explain why performance improves. This paper begins by exploring lossless and lossy communication, highlighting that the inclusion of data distortion distinguishes semantic communication from classical communication. It breaks the conditions for the separation theorem to hold and explains why the amount of data transferred by semantic communication is less. Therefor, employing JSCC becomes imperative for achieving optimal semantic communication. Moreover, a Variational Source-Channel Coding (VSCC) method is proposed for constructing semantic communication systems based on data distortion theory, integrating variational inference and channel characteristics. Using a deep learning network, we develop a semantic communication system employing the VSCC method and demonstrate its capability for semantic transmission. We also establish semantic communication systems of equivalent complexity employing the AE method and the VAE method. Experimental results reveal that the VSCC model offers superior interpretability compared to AE model, as it clearly captures the semantic features of the transmitted data, represented as the variance of latent variables in our experiments. In addition, VSCC model exhibits superior semantic transmission capabilities compared to VAE model. At the same level of data distortion evaluated by PSNR, VSCC model exhibits stronger human interpretability, which can be partially assessed by SSIM.
@@ II-C Variational Source-Channel Coding for Channel Matching
Figure 4: Geometric interpretation of the Variational Source-Channel Coding (VSCC) method. The raw data $x$ is a single sample from the original data distribution. It can be jointly encoded by VSCC encoder and the channel into the latent variable $z$. Each feature distribution represented by $z$ corresponds to a cluster of $x$, forming distinct regions. By resampling different feature distributions, the original data distribution can be reconstructed.
@@
$\min\text{KL}\left(\mathcal{N}\left(\mu_{x},\sigma_{1}^{2}+\sigma_{2}^{2}\right)\|\mathcal{N}\left(0,\sigma_{2}^{2}+d\right)\right)$ (9)
$=$ $\min\frac{1}{2}\left(\log\frac{\sigma_{2}^{2}+d}{\sigma_{1}^{2}+\sigma_{2}^{2}}+\frac{\mu_{x}^{2}+\sigma_{1}^{2}+\sigma_{2}^{2}}{\sigma_{2}^{2}+d}-1\right).$
@@
$\mathcal{L}_{\text{VSCC}}=$ $\mathbb{E}_{x\sim p(x)}\left[\frac{1}{2}\left(\log\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}(x;\theta)}\right.\right.$ (10)
$\left.\left.+\frac{\sigma_{1}^{2}(x;\theta)+\mu(x;\theta)^{2}}{\sigma_{2}^{2}}\right)-\log q_{x\mid z}(x\mid z)\right].$
@@ V Conclusion and Future Work
This paper identifies three key challenges in semantic communication. To address these issues, we analyze the differences between lossless and lossy communication, highlighting that for lossy communication, the assumptions required for the joint source-channel coding theorem are broken, rendering the separation theorem invalid. In this context, JSCC emerges as the optimal transmission strategy. Further, we highlight that the primary distinction between semantic communication and classical communication lies in whether the data is lossy. On the one hand, based on research in semantic information, it is clear that while semantics inherently depend on the existence of data, not all data necessarily contribute to semantics. Therefore, the goal of semantic encoding is to eliminate redundant data and extract its semantic features. Consequently, it significantly reduces the amount of data that needs to be transmitted, which is a major source of its gain. On the other hand, the presence of a knowledge base and specific communication tasks in semantic communication further expands beyond the boundary conditions of CIT, allowing joint coding to achieve optimal transmission while further reducing the data transmission load, which constitutes another key advantage of semantic communication.
As semantic communication requires a joint coding method that permits data distortion, we propose an innovative VSCC method. By introducing latent variables, the channel is integrated into the joint coding process. VSCC encodes original data into a channel-dependent feature distribution, allowing data to be categorized by the encoded feature distributions. This method enables data distortion while preserving the original data’s distributional characteristics. Unlike AE methods, VSCC focuses on lossless transmission of the original data distribution rather than the data itself.
We constructed the JSCC model using ResNet Blocks and trained it with VSCC, VAE, and AE methods. Experiments show that VSCC model effectively extracts common features from images, outputting them as the variances of the feature distributions. These variances, representing data dispersion, align with lossy transmission. Further, it can be found that the variances of feature distributions are influenced by the training channel SNRs, and varying CMCs during training leads to different transmission performances. It demonstrates that the channel is integrated into the production of feature distributions in VSCC, becoming part of the joint encoder rather than an obstacle. The VSCC model outperforms the VAE model in data recovery but falls short of the AE model, which targets lossless transmission. While an objective semantic evaluation metric is lacking, the comparison of PSNR and SSIM suggests that VSCC method achieves a degree of semantic recovery, with transmitted data more closely aligning with human perception.
@@
In our experiments, the VSCC method holds potential for further improvement. In Section 2.3, we assume that the encoded vector $Y$ follows the distribution $\mathcal{N}(\mu_{1},\sigma_{1}^{2})$, where the mean $\mu_{1}$ and variance $\sigma_{1}^{2}$ are learned according to Eq. 10. So the variance is constrained by CMC. Nevertheless, in the actual implementation of the VSCC model, after the joint encoder processes the original image $x$, it outputs the mean $\mu_{x}$ and variance $\sigma_{x}^{2}$, which together constitute the encoded vector $Y$. The mean $\mu_{x}$ can be interpreted as a sample drawn from the distribution $\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ and the variance $\sigma_{x}^{2}$ should tend to $\sigma_{1}^{2}$. If the first testing method from Section 3.3 is employed, both the mean $\mu_{x}$ and the variance $\sigma_{x}^{2}$ are transmitted. The received latent variable $Z$ should include the mean $\mu_{z}$ and the variance $\sigma_{z}^{2}$ that follow the distributions $\mathcal{N}(\mu_{1},\sigma_{1}^{2}+\sigma_{2}^{2})$ and $\mathcal{N}(\sigma_{1}^{2},\sigma_{2}^{2})$, respectively. Thus, resampling should be conducted according to the distribution $\mathcal{N}(\mathcal{N}(\mu_{1},\sigma_{1}^{2}+\sigma_{2}^{2}),\mathcal{N}(\sigma_{1}^{2},\sigma_{2}^{2}))$, which introduces an unstable variance. There should be a variance estimation process to correct the mean. However, this issue is overlooked in the VSCC loss function Eq. 10, leading to a decline in performance. Conversely, if the second testing method is adopted, where only the mean $\mu_{x}$ following the distribution $\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ is transmitted, it aligns seamlessly with the derivation of Eq. 10. However, because the true variance, following the $\mathcal{N}(\sigma_{1}^{2},\sigma_{2}^{2})$, is replaced by a fixed variance constrained by CMC, there would also be performance degradation. Consequently, if future work incorporates techniques from Weighted VAE to apply weights for multi-sample latent variables $Z$, where the weights can reflect the variance estimation process, the loss function of the VSCC method could approximate the maximum a posteriori estimation more closely, thereby enhancing the recovery of both data distribution and the data itself. It could potentially enable the VSCC model to surpass the performance of the AE model in testing. Additionally, the VSCC method illustrates that variance, as an indicator of data dispersion, is crucial for adapting to the channel and extracting shared semantic features from different images in semantic communication. This highlights the importance of estimation for the true variance, which happens to be a key topic in diffusion model research. By incorporating diffusion models to enhance the extraction and recovery of feature distributions represented by latent variables, the performance of the VSCC method could see significant improvement. Furthermore, the VSCC model currently lags behind the AE model when evaluated using existing metrics. This discrepancy does not necessarily indicate the inferiority of the VSCC method compared to the AE method. Instead, it arises from the fact that current evaluation metrics primarily focus on data recovery, which fails to capture the causal relationships between data distribution features and semantics. To further advance the field of semantic communication, it is essential to develop a comprehensive set of semantic-level evaluation metrics.
@@ Acknowledgment
The authors would like to thank…

### v2

@@ Abstract
The rapid development of artificial intelligence (AI) brings new opportunities and challenges to classical communication. AI applications not only impose significant transmission pressure on current communication networks but also affect the way information is stored and utilized due to its potent nonlinear fitting capabilities. Consequently, semantic communication technology emerges as a pivotal bridge connecting AI with classical communication. The current semantic communication systems are generally modeled as an Auto-Encoder (AE). AE lacks a deep integration of AI principles with communication strategies due to its inability to effectively capture channel dynamics. This gap makes it difficult to justify the need for joint source-channel coding (JSCC) and to explain why performance improves. This paper begins by exploring lossless and lossy communication, highlighting that the inclusion of data distortion distinguishes semantic communication from classical communication. It breaks the conditions for the separation theorem to hold and explains why the amount of data transferred by semantic communication is less. Therefore, employing JSCC becomes imperative for achieving optimal semantic communication. Moreover, a Variational Source-Channel Coding (VSCC) method is proposed for constructing semantic communication systems based on data distortion theory, integrating variational inference and channel characteristics. Using a deep learning network, we develop a semantic communication system employing the VSCC method and demonstrate its capability for semantic transmission. We also establish semantic communication systems of equivalent complexity employing the AE method and the VAE method. Experimental results reveal that the VSCC model offers superior interpretability compared to AE model, as it clearly captures the semantic features of the transmitted data, represented as the variance of latent variables in our experiments. In addition, VSCC model exhibits superior semantic transmission capabilities compared to VAE model. At the same level of data distortion evaluated by PSNR, VSCC model exhibits stronger human interpretability, which can be partially assessed by SSIM.
@@ II-C Variational Source-Channel Coding for Channel Matching
Figure 4: Geometric interpretation of the Variational Source-Channel Coding (VSCC) method. The raw data $x$ is a single sample from the original data distribution. It can be encoded by VSCC encoder and the channel into the latent variable $z$. Each feature distribution represented by $z$ corresponds to a cluster of $x$, forming distinct regions. By resampling different feature distributions, the original data distribution can be reconstructed and represented by $\hat{x}$.
@@
$\min\text{KL}\left(\mathcal{N}\left(\mu_{1},\sigma_{1}^{2}+\sigma_{2}^{2}\right)\|\mathcal{N}\left(0,\sigma_{2}^{2}+d\right)\right)$ (9)
$=$ $\min\frac{1}{2}\left(\log\frac{\sigma_{2}^{2}+d}{\sigma_{1}^{2}+\sigma_{2}^{2}}+\frac{\mu_{1}^{2}+\sigma_{1}^{2}+\sigma_{2}^{2}}{\sigma_{2}^{2}+d}-1\right).$
@@
$\mathcal{L}_{\text{VSCC}}=\mathbb{E}_{x\sim p(x)}\bigg{[}\frac{1}{2}\big{(}\log\frac{\sigma_{2}^{2}+d}{\sigma_{1}^{2}(x;\theta)+\sigma_{2}^{2}}$ (10)
$+\frac{\mu_{1}^{2}(x;\theta)+\sigma_{1}^{2}(x;\theta)+\sigma_{2}^{2}}{\sigma_{2}^{2}+d}-1\big{)}-\log q_{x\mid z}(x\mid z)\bigg{]}.$
@@ V Conclusion and Future Work
This paper identifies three key challenges in semantic communication. To address these issues, we analyze the differences between lossless and lossy communication, highlighting that for lossy communication, the assumptions required for the joint source-channel coding theorem are broken, rendering the separation theorem invalid. In this context, JSCC emerges as the optimal transmission strategy. Further, we highlight that the primary distinction between semantic communication and classical communication lies in whether the data is lossy. On the one hand, based on research in semantic information, it is clear that while semantics inherently depend on the existence of data, not all data necessarily contribute to semantics. Consequently, the goal of semantic encoding is to eliminate redundant data and extract its semantic features. The semantic communication significantly reduces the amount of data that needs to be transmitted, which is a major source of its gain. On the other hand, the presence of a knowledge base and specific communication tasks in semantic communication further expands beyond the boundary conditions of CIT, allowing joint coding to achieve optimal transmission while further reducing the transmission data, which constitutes another key advantage of semantic communication.
As semantic communication requires a joint coding method that permits data distortion, we propose an innovative VSCC method. By introducing latent variables, the channel is integrated into the joint coding process. VSCC encodes original data into a channel-dependent feature distribution, allowing data to be categorized by the encoded feature distributions. This method enables data distortion while preserving the original data’s distributional characteristics. Unlike AE methods, VSCC focuses on lossless transmission of the original data distribution rather than the data itself. We constructed the JSCC model using ResNet Blocks and trained it with VSCC, VAE, and AE methods. Experiments show that VSCC model effectively extracts common features from images, outputting them as the variances of the feature distributions. These variances, representing data dispersion, align with lossy transmission. Further, it can be found that the variances of feature distributions are influenced by the training channel SNRs, and varying CMCs during training leads to different transmission performances. It demonstrates that the channel is integrated into the production of feature distributions in VSCC, becoming part of the joint encoder rather than an obstacle. The VSCC model outperforms the VAE model in data recovery but falls short of the AE model, which targets lossless transmission. While an objective semantic evaluation metric is lacking, the comparison of PSNR and SSIM suggests that VSCC method achieves a degree of semantic recovery, with transmitted data more closely aligning with human perception.
@@
In our experiments, the VSCC method holds potential for further improvement. In Section 2.3, we assume that the encoded vector $Y$ follows the distribution $\mathcal{N}(\mu_{1},\sigma_{1}^{2})$, where the mean $\mu_{1}$ and variance $\sigma_{1}^{2}$ are learned according to Eq. 10. So the variance is constrained by CMC. Nevertheless, in the actual implementation of the VSCC model, after the joint encoder processes the original image $x$, it outputs the mean $\mu_{x}$ and variance $\sigma_{x}^{2}$, which together constitute the encoded vector $Y$. The mean $\mu_{x}$ can be interpreted as a sample drawn from the distribution $\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ and the variance $\sigma_{x}^{2}$ should tend to $\sigma_{1}^{2}$. If the first testing method from Section 3.3 is employed, both the mean $\mu_{x}$ and the variance $\sigma_{x}^{2}$ are transmitted. The received latent variable $Z$ should include the mean $\mu_{z}$ and the variance $\sigma_{z}^{2}$ that follow the distributions $\mathcal{N}(\mu_{1},\sigma_{1}^{2}+\sigma_{2}^{2})$ and $\mathcal{N}(\sigma_{1}^{2},\sigma_{2}^{2})$, respectively. Thus, resampling should be conducted according to the distribution $\mathcal{N}(\mathcal{N}(\mu_{1},\sigma_{1}^{2}+\sigma_{2}^{2}),\mathcal{N}(\sigma_{1}^{2},\sigma_{2}^{2}))$, which introduces an unstable variance. There should be a variance estimation process to correct the mean. However, this issue is overlooked in the VSCC loss function Eq. 10, leading to a decline in performance. Conversely, if the second testing method is adopted, where only the mean $\mu_{x}$ following the distribution $\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ is transmitted, it aligns seamlessly with the derivation of Eq. 10. However, because the true variance, following the $\mathcal{N}(\sigma_{1}^{2},\sigma_{2}^{2})$, is replaced by a fixed variance constrained by CMC, there would also be performance degradation. Consequently, if future work incorporates techniques from Weighted VAE to apply weights for multi-sample latent variables $Z$, where the weights can include the variance estimation process, the loss function of the VSCC method could approximate the maximum a posteriori estimation more closely, thereby enhancing the recovery of both data distribution and the data itself. It could potentially enable the VSCC model to surpass the performance of the AE model in testing. Additionally, the VSCC method illustrates that variance, as an indicator of data dispersion, is crucial for adapting to the channel and extracting shared semantic features from different images in semantic communication. It highlights the importance of estimation for the true variance, which happens to be a key topic in diffusion model research. By incorporating diffusion models to enhance the extraction and recovery of feature distributions represented by latent variables, the performance of the VSCC method could see significant improvement. Furthermore, the VSCC model currently lags behind the AE model when evaluated using existing metrics. This discrepancy does not necessarily indicate the inferiority of the VSCC method compared to the AE method. Instead, it arises from the fact that current evaluation metrics primarily focus on data recovery, which fails to capture the causal relationships between data distribution features and semantics. To further advance the field of semantic communication, it is essential to develop a comprehensive set of semantic-level evaluation metrics.
@@ Acknowledgment
We would like to express our sincere gratitude to Chulong Liang, Wei Zhao and Liguang Li from the Algorithm Department of ZTE Corporation for their invaluable support and insights during this research. Their expertise and guidance significantly contributed to the development of this work.